OpenStudy (anonymous):

if cos theta =2/3 find 2sec^2 theta + tan^2theta-7

OpenStudy (jtvatsim):

You will need to use the following pieces of information: \[\sin^2\theta + \cos^2\theta = 1\] \[\tan \theta = \frac{\sin \theta}{\cos \theta}\] \[\sec \theta = \frac{1}{\cos \theta}\]

OpenStudy (jtvatsim):

since all of these things are either sin or cos, and we already have cos, we should find sin first using the first formula.

OpenStudy (jtvatsim):

see if you can make any progress with that idea. :)